A uniform solid sphere of radius r=0.490 m and mass m=14.5 kg turns clockwise about a vertical axis through its center (when viewed from above), at an angular speed of 3.20 rad / s. What is its vector angular momentum about this axis?

(Enter the magnitude in kg · m² / s.)
Answer #1

Solution) R = 0.490 m

m = 14.5 kg

w = 3.20 rad/s

Angular momentum , L = ?

L = (I)(w)

I = (2/5)(m)(R^2)

I is moment of inertia of solid sphere

L = (2/5)(14.5)(0.490^2)(3.20)

L = 4.45 kgm^2/s

Direction is downwards